### v3

# Discovery and Preliminary Characterization of a Third Interstellar Object: 3I/ATLAS

Abstract
We report initial observations aimed at the characterization of a third interstellar object. This object, 3I/ATLAS or C/2025 N1 (ATLAS), was discovered on 2025 July 1 UT and has an orbital eccentricity of $e\sim 6.1$, perihelion of $q\sim 1.36$ au, inclination of $\sim 175^{\circ}$, and hyperbolic velocity of $V_{\infty}\sim 58$ km s-1. We report deep stacked images obtained using the Canada-France-Hawaii Telescope and the Very Large Telescope that resolve a compact coma. Using images obtained from several smaller ground-based telescopes, we find minimal light curve variation for the object over a $\sim 4$ day time span. The visible/near-infrared spectral slope of the object is 17.1$\pm$0.2 %/100 nm, comparable to other interstellar objects and primitive solar system small bodies (comets and D-type asteroids). 3I/ATLAS will be observable through early September 2025, then unobservable by Earth-based observatories near perihelion due to low solar elongation. It will be observable again from the ground in late November 2025. Although this limitation unfortunately prohibits detailed Earth-based observations at perihelion when the activity of 3I/ATLAS is likely to peak, spacecraft at Mars could be used to make valuable observations at this time.
[…]
{CJK*}
UTF8gbsn
show]dzs@msu.edu
1 Introduction
The first two interstellar objects identified traversing the inner Solar System, 1I/‘Oumuamua (G. V. Williams et al., 2017) and 2I/Borisov (G. Borisov et al., 2019), were discovered in 2017 and 2019, respectively. It has been suggested that interstellar objects formed in protostellar disks (A. Fitzsimmons et al., 2024) or the cores of giant molecular clouds (C.-H. Hsieh et al., 2021). Although there is little hope of identifying the exact home system for a given interstellar object (T. Hallatt & P. Wiegert, 2020), they provide the best opportunity to directly measure the properties of small bodies that formed outside of our solar system.
These first interstellar objects displayed divergent properties. For one, 1I/‘Oumuamua displayed no visible activity (K. J. Meech et al., 2017; Q.-Z. Ye et al., 2017; D. Jewitt et al., 2017; D. E. Trilling et al., 2018), yet nongravitational acceleration was detected in its trajectory at the $\sim 30\sigma$ level (M. Micheli et al., 2018). Meanwhile, 2I/Borisov displayed clear outgassing and dust ejection activity (D. Jewitt & J. Luu, 2019; A. Fitzsimmons et al., 2019; Q. Ye et al., 2020a; A. J. McKay et al., 2020; P. Guzik et al., 2020; M.-T. Hui et al., 2020; Y. Kim et al., 2020; G. Cremonese et al., 2020; B. Yang et al., 2021). The excess velocity of 1I/‘Oumuamua and 2I/Borisov also differed significantly. These objects had velocities of $V_{\infty}\sim 26$ km s-1 and $V_{\infty}\sim 32$ km s-1 respectively, which approximately correspond to ages of $\sim 10^{2}$ and $\sim 10^{3}$ Myr (E. Mamajek, 2017; E. Gaidos et al., 2017; F. Feng & H. R. A. Jones, 2018; F. Almeida-Fernandes & H. J. Rocha-Pinto, 2018; T. Hallatt & P. Wiegert, 2020; C.-H. Hsieh et al., 2021). In addition, 1I/‘Oumuamua displayed brightness variations of $\sim 3.5$ magnitudes corresponding to an extreme oblate $6:6:1$ geometry (K. J. Meech et al., 2017; M. M. Knight et al., 2017; W. C. Fraser et al., 2018; M. J. S. Belton et al., 2018; S. Mashchenko, 2019; A. G. Taylor et al., 2023) and had a moderately red reflectance spectrum (K. J. Meech et al., 2017; A. Fitzsimmons et al., 2018; Q.-Z. Ye et al., 2017).
These divergent properties led to a variety of hypotheses regarding the provenance of the population. Although 2I/Borisov’s coma was found to contain volatile species typically seen in comets (C. Opitom et al., 2019; T. Kareta et al., 2020; H. W. Lin et al., 2020; M. T. Bannister et al., 2020; Z. Xing et al., 2020; S. Bagnulo et al., 2021; K. Aravind et al., 2021), it had a high enrichment of CO relative to H2O (D. Bodewits et al., 2020; M. A. Cordiner et al., 2020). These ratios differentiate its composition from most solar system comets, which are typically rich in H2O and contain CO between 1–15% relative to water (N. Biver et al., 2024).
1I/‘Oumuamua’s lack of visible activity despite its nongravitational acceleration has led to a variety of hypothesized origins. M. Micheli et al. (2018) noted that for radiation pressure to cause the nongravitational acceleration, the object must either have an exceptionally low density or an extreme geometry. Such a density is a possible byproduct of diffusion-limited aggregation formation processes in the outskirts of a protostellar disk (A. Moro-Martín, 2019). Somewhat counterintuitively, it has also been demonstrated that such hypothetical structures bound together by weak van der Waals forces could survive tidal disruption from the Solar gravity (E. G. Flekkøy et al., 2019). It was also hypothesized that such a fractal aggregate could form in the coma of an undiscovered parent interstellar comet (J. X. Luu et al., 2020).
Alternatively, 1I/‘Oumuamua could have been outgassing volatiles with low levels of dust production, rendering it photometrically inactive in all extant observations (M. Micheli et al., 2018; Z. Sekanina, 2019; D. Seligman & G. Laughlin, 2020; W. G. Levine et al., 2021; W. G. Levine & G. Laughlin, 2021; S. J. Desch & A. P. Jackson, 2021; A. P. Jackson & S. J. Desch, 2021; S. J. Desch & A. P. Jackson, 2022; J. B. Bergner & D. Z. Seligman, 2023). This argument has been recently bolstered; since the discovery of 1I/‘Oumuamua, a series of photometrically inactive near-Earth objects (NEOs) have been reported to have significant comet-like nongravitational accelerations (D. Farnocchia et al., 2023; D. Z. Seligman et al., 2023, 2024). These objects imply that 1I/‘Oumuamua-like nongravitational accelerations may be more common than previously thought (A. G. Taylor et al., 2024). Regardless, our understanding of interstellar objects is incomplete. See A. Fitzsimmons et al. (2024), D. Z. Seligman & A. Moro-Martín (2023), D. Jewitt & D. Z. Seligman (2023), and A. Moro-Martín (2022) for recent reviews on this topic.
In this paper, we report early observations of 3I/ATLAS , the third interstellar object to be discovered after 1I/‘Oumuamua and 2I/Borisov, to help inform and coordinate follow-up observations.
2 Discovery and hyperbolic orbit characterization
2.1 Discovery
3I/ATLAS was discovered through the robotic observing schedule from ATLAS Chile (J. L. Tonry et al., 2018a) on 2025 July 1 and given the internal designation A11pl3Z [^1] https://www.minorplanetcenter.net/mpec/K25/K25N12.html (see Figure 1). The discovery tracklet was immediately submitted to the Minor Planet Center (MPC). Follow-up observations were then conducted by ATLAS in Hawai‘i, Sutherland, and the Canary Islands, along with dozens of other observatories worldwide. The discovery was made by ATLAS largely because the object was located in the Galactic plane — a region typically avoided by more sensitive surveys, such as Pan-STARRS and the Catalina Sky Survey.
Figure 1: Cutout images from the first and fourth discovery observations of 3I/ATLAS from the ATLAS Chile, spanning approximately one hour. 3I/ATLAS is moving at 0.49 deg/day against the stellar background. The cardinal directions and direction of motion are indicated with arrows, and 3I/ATLAS is identified within the red circle. (a) Un-background subtracted image from 05:15:11 UT; (b) Un-background subtracted image from 06:20:31 UT.
Six hours after the initial detection by ATLAS on 2025 July 1, pre-discovery detections from 2025 June 28–29 were identified in Zwicky Transient Facility (ZTF; E. C. Bellm et al., 2019; M. J. Graham et al., 2019) survey data, extending the orbital arc from 3.3 hours to 3 days. The new 3-day arc suggested a strongly hyperbolic orbit and prompted speculation on community mailing lists that the object might be interstellar. Later during the early hours of 2025 July 2 UT, additional pre-discovery pairs and triplets were identified in ATLAS data from five days earlier by S. Deen and later refined and resubmitted to the MPC by the ATLAS team. More pre-discovery detections from ZTF, dating back to 2025 June 14222Another batch of ZTF pre-discovery detections, dating back to 2025 May 22, were later submitted and published in MPEC 2025-N51 (https://minorplanetcenter.net/mpec/K25/K25N51.html)., were also identified and submitted.
At the time of discovery in the ATLAS data, the object had a magnitude 17.7–17.9 in the $o$-band filter (see Figure 1), and was at a heliocentric distance of $r=4.51$ au and a geocentric distance of $\Delta=3.50$ au.
Figure 2: Heliocentric orbit (ECLIPJ2000) of 1I/‘Oumuamua, 2I/Borisov, 3I/ATLAS, Earth, Mars, and Jupiter. Black lines represent the orbital path of each object; gray lines represent when the interstellar object is below-ecliptic. We highlight the location of the Sun (yellow) and the orbital paths of Earth (blue), Mars (red), and Jupiter (green). Large filled markers represent where the respective colored planet was when the interstellar object was discovered. Small unfilled markers represent where the planet will be at the interstellar object’s closest approach. Arrows point along the motion direction. The vernal equinox is indicated to the right.
2.2 Orbit Characterization
As of the early hours of 2025 July 2, the initial discovery and pre-discovery arc, covering a total of about 18 days, was sufficient to confirm the large eccentricity and, consequently, hyperbolic nature of the object’s orbit. As soon as the unusual nature of the object became evident, a large number of follow-up observations were obtained on 2025 July 2 by various observatories, leading to additional orbit refinement.
We first acquired 31 unfiltered 163 s exposures starting at 06:34 UT on 2025 July 2 spanning a total of 93 minutes using the European Space Agency’s (ESA) 0.56-meter Test Bed Telescope (TBT; MPC code W57) at La Silla Observatory in Chile, which has a field of view of 2.5 $\times$ 2.5. The telescope is fully devoted to NEO survey and follow-up observations, and can be interrupted at any time for high-profile targets.
[…]
A successful DDT request to the European Southern Observatory’s (ESO) 8.2-meter Very Large Telescope (VLT), submitted on the same day and quickly approved, also allowed our team to obtain images of the object during good seeing on 2025 July 4. High-precision astrometry (better than ±$\pm 0\farcs 1$) was extracted from this dataset, using an extrapolation to zero aperture to ensure correction for possible asymmetries in the inner coma of the object.
Our best estimate of the heliocentric orbital elements of the object, at the time of submission of this paper, is listed in Table 1. In Figure 2, we show the orbit of 3I/ATLAS in comparison to previously discovered interstellar objects.
Table 1: Initial orbit of 3I/ATLAS, computed using astrometry from 2025 May 22 to 2025 July 6. Heliocentric orbital elements at the epoch of 2025 July 3.
[…]
Perihelion distance $q$ [au] 1.3558541 $\pm$ 0.0019855
Eccentricity $e$ 6.1329072 $\pm$ 0.0162018
Inclination $i$ [] 175.11239 $\pm$ 0.00089
Longitude of ascending node $\Omega$ [] 322.14485 $\pm$ 0.01806
Argument of perihelion $\omega$ [] 128.01038 $\pm$ 0.02340
Time of perihelion $T_{P}$ [MJD, TDB] 60977.51085 $\pm$ 0.04078
From the current heliocentric orbit, it is possible to infer the incoming trajectory of the object before it entered our Solar System. The eccentricity of the object’s orbit with respect to the Solar System barycenter, computed before interacting with our planetary system, can be extrapolated as $e_{b}=6.144\pm 0.016$. This value, together with an incoming pericenter distance of $q_{b}=(1.3611\pm 0.0020)$ au, results in an incoming velocity $v_{\infty}=(57.942\pm 0.049)$ km s-1, from an asymptote directed towards a Right Ascension of $\sim 295^{\circ}$ and a Declination of $\sim-19^{\circ}$, in the constellation of Sagittarius and not far from the Galactic Center.
3 Observational Analysis
3.1 Activity
Given the stark difference in visible activity levels between 1I/‘Oumuamua and 2I/Borisov (Section 1), a high priority for initial observational analysis for 3I/ATLAS was determining whether it was active or not. With this in mind, we obtained three 60 s non-sidereally guided gri-band images with the MegaCam wide-field mosaic imager (O. Boulade et al., 2003) on the 3.6-meter Canada-France-Hawaii Telescope (CFHT, MPC code T14) on 2025 July 2 to search for faint cometary activity. The highest-quality image had typical stellar FHWMs of $0\farcs 72\pm 0\farcs 05$ measured perpendicular to the direction of trailing, while the object had a FWHM of $1\farcs 29\pm 0\farcs 02$. The magnitude in a $5\farcs 0$ radius aperture was 17.2 in the Gaia DR2 G band after three background stars were masked, although the field was crowded, so we consider this measurement to be unreliable. Figure 3a shows the stacked composite image of these data in which faint activity is visible.
On 2025 July 4, fifteen 10 s R-band images were acquired using the FORS2 instrument on the 8.2-meter Unit Telescope 1 of the European Southern Observatory Very Large Telescope (VLT) on Cerro Paranal, Chile. The measured stellar image quality was $0\farcs 6$ FWHM. However, the object had a FWHM of $1\farcs 49\times 1\farcs 29$ with the asymmetric elongation pointing to position angle $277^{\circ}$ from North to East, where the coma can be seen extending out to $3\farcs 5$. A composite image of these data is shown in Figure 3b. The photometric profile of the object (Figure 4) shows a brightness excess extending over $4\farcs 0$ compared to that of field stars.
Figure 3: (a) Stacked gri-band image cut-out from CFHT on 2025 July 2 and (b) R-band composite image from VLT on 2025 July 4, both showing faint activity. Background objects were removed from the VLT stack. Arrows indicate the directions of North (N), East (E), the anti-solar vector as projected on the sky (direct-product-$\odot$- ⊙), and the negative heliocentric velocity vector as projected on the sky ($-v$).
Figure 4: Photometric profile of the VLT image (Figure 3a). The individual pixels are represented by black dots; average in annuli by blue circles. The red line is the average stellar profile measured on 10 nearby well-exposed stars.
Figure 5: Absolute magnitudes computed from ZTF precovery data published in MPEC 2025-N51 from 2025 May 22 to June 18 (black filled dots). The median $H_{V}$ and standard deviation of data are shown by a dashed horizontal line and dotted horizontal lines, respectively. No uncertainties were reported for these original data, and so plotted uncertainties are estimated based on the standard deviation of photometric points measured between 2025 May 22 and June 14. Corresponding pre-perihelion $H_{V}$ for 2I/Borisov derived from Q. Ye et al. (2020b) are plotted on the same heliocentric distance ($r$) scale (open circles) for comparison.
As noted in Section 2, $g^{\prime}$- and $r^{\prime}$-band precovery data from ZTF were published in MPEC 2025-N51 following these observations. This included photometry dating back to These data mostly consisted of single $g^{\prime}$- or $r^{\prime}$-band detections on each night, but included pairs of $g^{\prime}$ and $r^{\prime}$-band detections on 2025 May 22 and June 18. We compute inferred $V$-band absolute magnitude from these data assuming a standard asteroidal phase function (E. Bowell et al., 1989) assuming $G=0.15$ and solar colors (K. Jordi et al., 2006; J. Holmberg et al., 2006). We use this to present a secular light curve in Figure 5. The computed absolute magnitudes are relatively constant between 2025 May 22 and June 14 (during which the object ranged between heliocentric distances of $r=5.8$ au to $r=5.1$ au). The secular light curve has a median value of $H_{V}\sim 12.4$ and a standard deviation of $\sim$0.1 mag.
The photometry from 2025 June 18 (the last night plotted in Figure 5 when the object was at $r=4.9$ au) corresponds to an absolute magnitude outside that standard deviation at $H_{V}=12.1\pm 0.1$. For comparison, $r^{\prime}$-band light curve data obtained from LCO-network telescopes (discussed in more detail in Section 3.2) show an approximate average $r^{\prime}$-band apparent magnitude of $m_{r}\sim 17.9$, corresponding to $H_{V}=11.9$, indicating that the faint coma visible in CFHT and VLT data obtained on 2025 July 2 and 4 could be responsible for about 0.5 mag of excess brightening of the object since early June.
We note that this finding does not exclude the presence of coma prior to mid-June, but indicates that any coma that may have been present was relatively steady-state from late May to early June, and then increased in mid-June. For reference, $H_{V}=12.4\pm 0.1$ corresponds to a radius of $\sim(10\pm 1)$ km for a geometric albedo of $p_{V}=0.05$ typical of comets, but this is an upper limit given that we cannot rule out activity during the observations used to derive this result.
[…]
3.2 Rotation
Table 2: Light Curve Observations Summary
MPC Observation Pixel Photometry Aperture
Facility/Instrument Site Code Dates (UT) Filter(s) Scale Method Radiusa
ATLAS (Haleakala) W68 2025 July 4 $o$ $1\farcs 86$ PSF n/a
ATLAS (Chile) T05 2025 July 4 $o$ $1\farcs 86$ PSF n/a
Faulkes Telescope North F65 2025 July 2, 4 $r^{\prime}$ $0\farcs 28$ aperture $2\farcs 2$
Faulkes Telescope South E10 2025 July 2, 4 $r^{\prime}$ $0\farcs 28$ aperture $2\farcs 2$
LCO 0.35m (Haleakala #1) T03 2025 July 2, 4 $r^{\prime}$ $0\farcs 74$ aperture $3\farcs 0$
LCO 0.35m (Haleakala #2) T04 2025 July 2, 4 $r^{\prime}$ $0\farcs 74$ aperture $3\farcs 0$
Telescope Joan Oró C65 2025 July 4-5 $r^{\prime}$ $0\farcs 72$ aperture $2\farcs 2$
TRAPPIST-North Z53 2025 July 2-4 $R_{c}$ $1\farcs 20$ aperture $4\farcs 8$
TRAPPIST-South I40 2025 July 2-4 $R_{c}$ $1\farcs 28$ aperture $5\farcs 1$
a Photometry aperture size if aperture photometry is used; otherwise, this field is not applicable (n/a).
[…]
Photometric observations were obtained on the nights of 2025 July 2–4 with the TRAPPIST telescopes (E. Jehin et al., 2011). TRAPPIST-North (TN, MPC code Z53) is located at the Oukaimeden observatory in Morocco, and TRAPPIST-South (TS, MPC code I40) is located at the ESO La Silla Observatory in Chile. Both telescopes are 0.6-meter Ritchey-Chretien telescopes operating at f/8. The TN camera is an Andor IKONL BEX2 DD camera providing a $20^{\prime}$ field of view and pixel scale of . TS is equipped with a FLI ProLine 3041-BB CCD camera with a $22^{\prime}$ field of view and a pixel scale of . Johnson-Cousin $R_{c}$ filters were used with individual exposure times of 180 s .
[…]
Compiled light curve results are shown in Figure 6. .
No periodicity is immediately apparent within any of the datasets, indicating little brightness variation during the observation period This is in notable contrast to 1I/‘Oumuamua, which displayed a far more extreme light curve range of $\sim$ 3.5 magnitudes (see Section 1). It should be noted, however, that the detection of activity and coma (Section 3.1) means that any photometric variations caused by a rotating nucleus will be at least partially suppressed by the relatively more steady state coma surrounding it.
3.3 Colors
[…]
The colors of 3I/ATLAS as seen in Figure 7 are relatively linear (e.g., without obvious absorption features or spectral curvature) and significantly redder than those of the Sun. This slope of this reflectance spectrum is approximately 18 %/100 nm as derived from the full range of the available color data. It is similar, though somewhat redder, to that retrieved for 1I/‘Oumuamua (10$\pm$6 %/100 nm, Q.-Z. Ye et al., 2017) and 2I/Borisov (12$\pm$1 %/100 nm J. de León et al., 2020), both of which are slightly redder than the D-type asteroids (F. E. DeMeo et al., 2009). However, it is not as red as seen in some outer Solar System objects (e.g. Pholus, R. P. Binzel, 1992). Some of the reflected light must be from the coma because 3I/ATLAS is weakly active. This is corroborated by the fact that photometry of $g^{\prime}$- and $r^{\prime}$-band ZTF precovery data (Section 3.1) obtained when the object appeared to be less active, indicated that the $g^{\prime}-r^{\prime}$ colors on those nights were close to solar (within large uncertainties). Specifically, the measurements were $(g^{\prime}-r^{\prime})=0.42\pm 0.14$ on 2025 May 22 and $(g^{\prime}-r^{\prime})=0.44\pm 0.14$ on June 18, compared to $(g^{\prime}-r^{\prime})_{\odot}=0.45\pm 0.02$ (J. Holmberg et al., 2006). This suggests that the much redder color measured from later data could be heavily affected by ejected dust. The spectral slope measured here is slightly higher than the average for cometary dust (see modeling and discussions of typical Solar System comets in S. Protopapa et al. 2018; T. Kareta et al. 2023, as well as Figure 3 of L. Kolokolova et al. 2024), but not significantly so. 2I/Borisov’s coma was similarly red (see, e.g., J. de León et al. 2020), but the object was also significantly more active at the time of the color observations.
3.4 Spectrum
The spectrum shown in Figure 7 was obtained on 2025 July 4 using the SNIFS instrument (SuperNova Integral Field Spectrograph; B. Lantz et al., 2004) on the UH 2.2-meter telescope at Maunakea by the Spectroscopic Classification of Astronomical Transients (SCAT; M. A. Tucker et al., 2022) team. Data reduction followed the procedures described in M. A. Tucker et al. (2022). SNIFS contains two channels split by a dichroic mirror, with a blue channel covering 0.34–0.51 $\mu$m, and a red channel covering 0.51–1.0 $\mu$m, with spectral resolutions of 5 Å and 7 Å, respectively. The fast reduction pipeline used here does not include the full dichroic correction; therefore, data in this region were excluded from the final spectrum. In addition to 3I/ATLAS, the solar analog HD 165290 was observed. The final calibrated reflectance spectrum obtained by dividing the target spectrum by the normalized spectrum of the solar analog is shown in Figure 7.
The normalized reflection spectrum shows no obvious absorption features or gas emissions and yields a spectral slope of 17.1$\pm$0.2 %/100 nm . This value was determined with a linear fitting routine with bootstrapping to estimate uncertainties. This spectral slope confirms the moderately red slope observed with the MuSCAT3 images, further solidifying a color similarity to both D-types and 2I/Borisov.
4 Discussion
Given that three interstellar objects have now been discovered, it is worthwhile to use their measured physical properties and discovery circumstances to estimate the size of the total population in the Solar System.
The volume within which ATLAS can detect a moving object depends strongly on intrinsic brightness ($H_{V}$) and location in the Solar System, where sensitivity, illumination, distance, phase function, and trailing losses are critical components. Using an ATLAS detection model that assumes a $G=0.15$ phase function, a speed of 50 km s-1 for trailing losses, a weather and moon averaged limiting $m_{lim}\simeq 19$, and 50 km s-1 speed for volume crossing, Figure 8 shows the volume in which ATLAS can detect an object.
[…]
This curve of detectability volume divided by crossing time has a slope of $\log(V/dt)\propto-0.2H_{V}$ for $H_{V}<15$ and a slope of $\log(V/dt)\propto-0.4H_{V}$ for $H_{V}>15$. However, the crossing time for $H>16$ is less than 0.1 yr which means that the actual detectability is certainly less than $V/dt$ from Figure 8 because of weather and other survey cadence effects. The ratio of detectability volume and crossing time is approximately the product of detection cross section and velocity, hence the ratio of detection rate and local density. Evaluating Figure 8 at $H_{V}\sim 12.5$ for 3I/ATLAS gives $600$ au3 yr-1, and a local density of $\sim 3\times 10^{-4}$ au-3 for a detection rate of $\sim$ 0.2 yr-1.
The $G=0.15$ phase function used for Figure 8 may be inappropriate for an object such as 3I/ATLAS with an unknown amount of coma, and unknown and potentially unfamiliar surface scattering properties, which could cause this density estimate to be significantly incorrect. Obtaining stronger constraints on 3I/ATLAS’s nucleus size, such as with high-resolution imaging by space telescopes or adaptive optics-equipped ground-based telescopes, or by searching for occultation observation opportunities, should therefore be considered a high priority.
A. Do et al. (2018) estimated an interstellar density of objects like 1I/‘Oumuamua at 0.2 au-3 ($H_{V}\sim 23$) and noted that the size distribution must be steeper than $n(d>D)\propto D^{-3}$ to avoid a logarithmic divergence in total mass, i.e. differential $\log n(d)$ with slope greater than $0.8H_{V}$ (see Fig. 3 of A. Do et al., 2018). This criterion applied to the ATLAS detectability curve would cause a strong peak around $H\sim 16$, not two detections at $H\sim 13$, however. As such, the detections of 2I/Borisov and 3I/ATLAS, both of which appear, or appeared, relatively large, are somewhat unexpected given that more numerous smaller interstellar objects have not also been discovered over the same time period. Clearly what is needed is more interstellar object discoveries to improve population statistics, which is hopefully exactly what the Rubin Observatory Legacy Survey of Space and Time (LSST; R. L. Jones et al., 2009) will deliver in the coming years.
We are certain to learn more about 3I/ATLAS in the coming weeks and months as it experiences continued and increasing heating for perhaps the first time during its passage past our Sun. Comprehensive and collaborative investigations of 3I/ATLAS based off the lessons learned from the 1I/‘Oumuamua and 2I/Borisov campaigns are poised to significantly expand our knowledge of the interstellar object population, and as such, additional observations of 3I/ATLAS are highly encouraged. Further photometry, spectroscopy, or polarimetry could constrain the rotation, activity, dust size-frequency distribution, coma composition, and nongravitational acceleration of the object. The perihelion of 3I/ATLAS will not be easily observable from Earth-based observatories, as the object will be on the opposite side of the Sun and at a low solar elongation angle. However, the object will approach within 0.19 au of Mars, and we encourage nearby spacecraft (e.g., Mars Reconnaissance Orbiter, Trace Gas Orbiter, MAVEN, Tianwen-1, Hope, Jupiter Icy Moons Explorer) equipped with visible, UV, and IR spectrographs and cameras to attempt to capture data on this object’s closest approach.
5 Data and Software Availability
Data behind many of the figures in this work are available at https://github.com/3I-ATLAS/discovery-paper.
6 acknowledgments
D.Z.S. is supported by an NSF Astronomy and Astrophysics Postdoctoral Fellowship under award AST-2303553. This research award is partially funded by a generous gift of Charles Simonyi to the NSF Division of Astronomical Sciences. The award is made in recognition of significant contributions to Rubin Observatory’s Legacy Survey of Space and Time. D.F. conducted this research at the Jet Propulsion Laboratory, California Institute of Technology, under a contract with the National Aeronautics and Space Administration (80NM0018D0004). T.S-R. acknowledges funding from Ministerio de Ciencia e Innovación (Spanish Government), PGC2021, PID2021-125883NB-C21. This work was (partially) supported by the Spanish MICIN/AEI/10.13039/501100011033 and by “ERDF A way of making Europe” by the “European Union” through grant PID2021-122842OB-C21, and the Institute of Cosmos Sciences University of Barcelona (ICCUB, Unidad de Excelencia ‘María de Maeztu’) through grant CEX2019-000918-M. A.D.F. acknowledges funding from NASA through the NASA Hubble Fellowship grant HST-HF2-51530.001-A awarded by STScI. K.J.M., J.W., and A.H. acknowledge support from the Simons Foundation through SFI-PD-Pivot Mentor-00009672. A.G.T. acknowledges support from the Fannie and John Hertz Foundation and the University of Michigan’s Rackham Merit Fellowship Program. D.M. acknowledges support by the Science Fund of the Republic of Serbia, GRANT No. 7453, Demystifying enigmatic visitors of the near-Earth region (ENIGMA). E.P.-A. acknowledges support by the Italian Space Agency within the LUMIO project (ASI-PoliMi agreement n. 2024-6-HH.0). B.J.S, K.H., and W.B.H. acknowledges support from NSF (grants AST-2407205) and NASA (grants HST-GO-17087, 80NSSC24K0521, 80NSSC24K0490, 80NSSC23K1431). W.B.H acknowledges support from the National Science Foundation Graduate Research Fellowship Program under Grant Nos. 1842402 and 2236415.
This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant Nos. 1842402 and 2236415. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. K.E.M. acknowledges support from the NASA ROSES DDAP program funded through NASA Goddard Space Flight Center.
This work has made use of data from the Asteroid Terrestrial-impact Last Alert System (ATLAS) project. ATLAS is primarily funded to search for near-Earth asteroids through NASA grants NN12AR55G, 80NSSC18K0284, and 80NSSC18K1575; byproducts of the NEO search include images and catalogs from the survey area. The ATLAS science products have been made possible through the contributions of the University of Hawaii Institute for Astronomy, the Queen’s University Belfast, the Space Telescope Science Institute, the South African Astronomical Observatory (SAAO), and the Millennium Institute of Astrophysics (MAS), Chile.
This work also includes observations obtained with MegaPrime/MegaCam, a joint project of CFHT and CEA/DAPNIA, at the Canada-France-Hawaii Telescope (CFHT) which is operated by the National Research Council (NRC) of Canada, the Institut National des Science de l’Univers of the Centre National de la Recherche Scientifique (CNRS) of France, and the University of Hawaii. The observations at the Canada-France-Hawaii Telescope were performed with care and respect from the summit of Maunakea which is a significant cultural and historic site.
This paper is based on observations made with the MuSCAT3 instrument, developed by Astrobiology Center and under financial supports by JSPS KAKENHI (JP18H05439) and JST PRESTO (JPMJPR1775), at Faulkes Telescope North (FTN) on Maui, Hawaii, operated by the Las Cumbres Observatory. Coordination of some observations conducted using LCO network telescopes was carried out as part of the LCO Outbursting Objects Key project (LOOK; T. Lister et al., 2022).
The Joan Oró Telescope (TJO) at the Montsec Observatory (OdM) is owned by the Catalan Government and operated by the Institute of Space Studies of Catalonia (IEEC).
This research has made use of data and/or services provided by the International Astronomical Union’s Minor Planet Center.